OpenStudy (aaronandyson):

what is the abs val of -4- sq root of 2 i

OpenStudy (aaronandyson):

-4-√2i

sam (.sam.):

\[|-4-\sqrt{2}i| \\ \\ = |-1||4+ \sqrt{2}i|\] Since \(-1\leq 0\), |-1| = 1 \[=|4+\sqrt{2}i|\] Now use the formula \[|z|=|a+bi| = \sqrt{a^2+b^2}\]

OpenStudy (aaronandyson):

3√2*i?

sam (.sam.):

Yep, and without the imaginary unit

OpenStudy (aaronandyson):

So that 3√2

sam (.sam.):

Mhm

myininaya (myininaya):

just for fun @AaronAndyson you could have pretended the question was what is distance from (-4,-sqrt(2)) to (0,0) and referred to your old distance formula \[d=\sqrt{(-4-0)^2+(-\sqrt{2}-0)^2} \\ d=\sqrt{(-4)^2+(-\sqrt{2})^2}\] and then continue simplifying it from here I hope you can see that the questions are equivalent and the formulas sam and I use are equivalent since we are finding a distance between a point given and the origin point

OpenStudy (aaronandyson):

Thanks for the suggestion @myininaya

OpenStudy (agent0smith):

For future reference, the absolute value of a complex number, refers to its magnitude. You don't need to factor out a -1 as Sam did. All you need is the formula that was given at the end of the post \[\large |z|=|a+bi| = \sqrt{a^2+b^2}\]